# iLRM: An Iterative Large 3D Reconstruction Model

### v1

@@ Abstract
Feed-forward 3D modeling has emerged as a promising approach for rapid and high-quality 3D reconstruction. In particular, directly generating explicit 3D representations, such as 3D Gaussian splatting, has attracted significant attention due to its fast and high-quality rendering, as well as numerous applications. However, many state-of-the-art methods, primarily based on transformer architectures, suffer from severe scalability issues because they rely on full attention across image tokens from multiple input views, resulting in prohibitive computational costs as the number of views or image resolution increases. Toward a scalable and efficient feed-forward 3D reconstruction, we introduce an iterative Large 3D Reconstruction Model (iLRM) that generates 3D Gaussian representations through an iterative refinement mechanism, guided by three core principles: (1) decoupling the scene representation from input-view images to enable compact 3D representations; (2) decomposing fully-attentional multi-view interactions into a two-stage attention scheme to reduce computational costs; and (3) injecting high-resolution information at every layer to achieve high-fidelity reconstruction. Experimental results on widely used datasets, such as RE10K and DL3DV, demonstrate that iLRM outperforms existing methods in both reconstruction quality and speed. Notably, iLRM exhibits superior scalability, delivering significantly higher reconstruction quality under comparable computational cost by efficiently leveraging a larger number of input views.
@@ 1 Introduction
Since the recent success of 3D Gaussian Splatting (3D-GS) kerbl20233dgs , significant progress has been made in leveraging this 3D representation for building generalizable feed-forward 3D reconstruction models charatan2024pixelsplat ; tang2025lgm ; xu2024grm ; chen2025mvsplat ; chen2024mvsplat360 ; xu2024depthsplat ; zhang2025gs-lrm . These methods typically train large neural networks to transform multi-view input images into feature representations, then regress Gaussian attributes. In contrast to per-scene 3D-GS optimization approaches kerbl20233dgs ; mallick2024taming ; lu2024turbo-gs ; fang2024mini-splat , these feed-forward models can reconstruct 3D scenes in a single forward pass, offering near real-time performance. Moreover, the prior knowledge learned from large-scale datasets zhou2018re10k ; ling2024dl3dv ; deitke2023objaverse ; deitke2024objaverse-xl allows them to effectively generalize to unseen scenes. While their reconstruction quality often lags behind that of per-scene optimization methods, fast reconstruction speed and generalization capability mark a promising step toward the long-standing goal of achieving accurate and real-time 3D scene reconstruction.
@@
Among the promising approaches, pixel-aligned Gaussian models charatan2024pixelsplat ; szymanowicz2024splatter ; zheng2024gps have emerged as the de facto standard, leveraging decades of advances in network architectures developed for numerous image-based tasks. While these models have proven effective, they also exhibit certain limitations. In particular, since they generate per-pixel Gaussians directly from the input images, the image resolution determines the number of Gaussians produced, which can lead to an excessive number of redundant Gaussians. For example, given input images at 1K resolution across 200 viewpoints (a scale comparable to the bicycle scene in the mip-NeRF 360 dataset barron2022mip ), these methods would produce 200 million Gaussians, despite prior studies lee2024compact ; fang2024mini-splat ; fan2024lightgaussian ; chen2024hac ; lee2025optimized demonstrating that such scenes can be efficiently represented with around 0.5 million Gaussians. To mitigate this issue, several techniques have been proposed, such as Gaussian regularization ziwen2024long-lrm and feature fusion wang2025freesplat . Alternatively, the network architecture can also be designed to generate fewer Gaussians, for example by downsampling the output resolutions. However, these strategies still require processing high-resolution multi-view images and therefore do not address another fundamental limitation of these models: high computational and memory demands.
@@
A significant portion of computational and memory overhead arises from modeling interactions across multiple input views in feed-forward 3D reconstruction models. For instance, GS-LRM zhang2025gs-lrm performs full attention over all image tokens from every input view, leading to a quadratic increase in complexity with respect to both the number of views and image resolution. MVSplat chen2025mvsplat and DepthSplat xu2024depthsplat construct and process dense cost volumes for each view, further contributing to the overall computational demands. While one might attempt to alleviate this burden by reducing the input image resolution or using a sparser set of views, such strategies risk discarding essential geometric and appearance information required for accurate reconstruction.
@@
Beyond the computational complexity and the inefficiency of the generated representations, we also question whether the prevailing formulation, casting 3D reconstruction as a sequence-to-sequence problem that maps entire sets of image tokens to dense, pixel-aligned Gaussians, is fundamentally well-suited to the nature of the task. While this formulation has achieved impressive results (even without explicit 3D inductive biases zhang2025gs-lrm ; ye2024no ), it remains primarily a one-shot 3D scene generation process. In contrast, the recent optimization-based methods kerbl20233dgs ; mallick2024taming follow a fundamentally different strategy: they treat reconstruction as an iterative refinement process, where each iteration involves rendering the current scene estimate, measuring reconstruction error, and updating the representation accordingly. This loop enables the model to progressively capture finer geometric and appearance details while ensuring strong 3D consistency. The success of these methods suggests that high-quality 3D reconstruction may benefit not only from expressive representations but also from feedback-driven iterative refinement, a trait largely absent in existing feed-forward 3D models.
@@
We have comprehensively evaluated the proposed method on the large-scale datasets, RealEstate10K zhou2018re10k and DL3DV ling2024dl3dv . The experimental results demonstrate that iLRM achieves superior rendering quality while significantly reducing computational and memory overhead compared to recently proposed feed-forward Gaussian models. For instance, on the RealEstate10K dataset, iLRM improved the PSNR by 3$\sim$ 3∼ 3 dB over the state-of-the-art methods, such as GS-LRM zhang2025gs-lrm and DepthSplat xu2024depthsplat , by leveraging more views (8 views vs. 2 views) with only less than half of the computation time (0.028s vs. 0.065s). We further demonstrated that our method can generalize well across diverse scenes and viewpoint configurations, making it suitable for practical, large-scale 3D reconstruction tasks.
2 Related Works
2.1 Feed-forward 3D Gaussian Splatting
Feed-forward 3D Gaussian Splatting charatan2024pixelsplat ; chen2025mvsplat ; xu2024depthsplat ; nam2024generative ; zhang2025gs-lrm ; szymanowicz2024splatter ; tang2025lgm ; xu2024grm capitalizes on robust priors learned from extensive datasets to estimate Gaussian primitive parameters and synthesize novel view images using sparse input data. PixelSplat charatan2024pixelsplat and LatentSplat wewer2024latentsplat predict Gaussians from image features using an epipolar line sampling method to enhance geometric accuracy, while MVSplat chen2025mvsplat and MVSGaussian liu2024mvsgaussian construct cost volumes through a plane-sweep stereo approach. In a further development, Flash3D szymanowicz2024flash3d and DepthSplat xu2024depthsplat introduce a pre-trained depth estimation model piccinelli2024unidepth ; depthanything , which improves the robustness of the spatial positions of 3D Gaussians. In contrast, GS-LRM zhang2025gs-lrm and Long-LRM ziwen2024long-lrm minimize reliance on explicit 3D priors by leveraging large-scale data-driven priors and pre-trained 2D foundation models caron2021dino-v1 ; oquab2023dino-v2 ; depthanything .
Iterative refinements. Our work is also closely related to recent methods that adopt iterative refinement strategies, such as G3R chen2024g3r and Gen-Den nam2024generative . Both utilize actual gradients to update their representations more precisely. While promising, these approaches require additional computational burden for rendering multiple images per training iteration, and relying solely on gradients may risk overlooking valuable information present in the raw input images. Nonetheless, exploring how to incorporate gradient information remains an interesting direction for future work.
2.2 3D Representations from Embeddings
Inspired by previous generative approaches goodfellow2014generative ; karras2019style ; chan2022efficient , recent works hong2023lrm ; chen2024lara ; flynn2024quark have investigated the synthesis of 3D representations directly from learnable embeddings, guided by input image supervision. This paradigm leverages the expressive capacity of latent spaces to encode rich geometric priors, which act as structural templates that guide the reconstruction process. Such approaches offer notable flexibility, allowing rendering from arbitrary viewpoints and adaptation to varying space scales and camera poses. However, both LRM hong2023lrm and Lara chen2024lara are limited to object-centric representations, restricting scalability to complex scenes involving multiple objects or large spatial layouts. The recently proposed Quark flynn2024quark also utilizes learnable embeddings to fuse visual cues from multiple images, demonstrating compelling results, but its representation is confined to the target view xu2024murf ; liu2024mvsgaussian ; jin2024lvsm , lacking an explicit and persistent 3D reconstruction.
@@ 3 Method
In this section, we provide a comprehensive explanation of our method, iLRM. We begin with the motivation and problem statement in Sec. 3.1, followed by a detailed presentation of the architectural design in Sec. 3.2, and finally training objectives in Sec. 3.3.
@@ 3.1 Motivation and Problem Statement
Existing generalizable 3D Gaussian reconstruction methods process multi-view images in an end-to-end fashion, often employing techniques such as epipolar line sampling charatan2024pixelsplat , plane-sweep stereo chen2025mvsplat ; chen2024mvsplat360 ; xu2024depthsplat , or full-resolution attention zhang2025gs-lrm ; xu2024grm ; ziwen2024long-lrm to enforce multi-view consistency. While effective, these strategies introduce significant computational and memory overhead, limiting their scalability.
@@
We start with $N$ multi-view images $\{I_{i}\}_{i=1}^{N}$ and their corresponding camera poses $\{C_{i}\}_{i=1}^{N}$. Based on this setup, our goal is to train a model $f_{\theta}$ that maps a set of camera viewpoints to 3D Gaussians, leveraging the associated multi-view images as visual cues throughout the reconstruction pipeline. More formally,
@@ 3.2 Architectural Design
Figure 2: The proposed scalable architectural designs by decoupling viewpoint and image tokens, and modeling the global interactions via cross- and self-attentions ($N$: # views, $h=H/p,w=W/p$).
Viewpoint tokenization. Following previous works tang2025lgm ; zhang2025gs-lrm ; jin2024lvsm , we employ a Plücker ray embedding for each input view using the camera poses. Specifically, given the intrinsic, rotation, and translation, we construct the Plücker ray embeddings for each viewpoint. We then divide these viewpoint embeddings into non-overlapping patches of size $p\times p$, and reshape each patch into a 1D vector, resulting in a tensor of shape $H^{v}W^{v}/p^{2}\times 6p^{2}$. Then, we encode it using a single linear layer to produce viewpoint tokens, $V_{i}^{(0)}\in\mathbb{R}^{H^{v}W^{v}/p^{2}\times d}$. Plücker coordinates inherently capture spatial variations across pixels and views, allowing them to effectively differentiate between patches. As a result, we do not utilize additional positional embeddings.
@@
Scalable multi-view context modeling. Fig. 2-(a) shows the typical feed-forward 3D methods zhang2025gs-lrm ; chen2025mvsplat ; xu2024depthsplat using transformer architecture, which perform full attention across multi-view images, incurring a quadratic increase in computational cost with respect to both the number of views and the image resolution. Fig. 2-(b) depicts our decoupling approach. Thanks to the decoupling technique, we can reduce the viewpoint resolution while still leveraging high-resolution multi-view images (Fig. 2-(c)). We further decrease the computation cost by two-stage multi-view context modeling, per-view cross-attention and viewpoint self-attention (Fig. 2-(d)). For example, given 16 input images with a resolution of 256 $\times$ 256 and a patch size of 8, the relative computational cost follows the ratio (1:1:0.25:0.08, Fig. 2-(a):(b):(c):(d)), highlighting that our proposed method can accommodate more input views with significantly less computational burden.
@@
Figure 4: Various mini-batch cross-attention schemes.
@@ 3.3 Training Objectives
After generating 3D Gaussians from viewpoint tokens, we rasterize them to the target viewpoint to obtain rendered images, $\hat{I}_{t}$. These rendered images are then supervised using ground-truth images, $I_{t}$, through MSE loss and perceptual loss chen2017photographic ; li2020crowdsampling using a pretrained VGG network simonyan2014very . For each training scene, our training loss function is as follows.
@@ 4.1 Datasets
We train our model on two large-scale datasets: RealEstate10K (RE10K) zhou2018re10k , DL3DV ling2024dl3dv , and evaluate it on three datasets, including ACID liu2021acid , which is used only for evaluation. For RE10k, we use 67,477 training and 7,289 testing scenes; for DL3DV, 9,568 training and 140 testing scenes, consistent with previous works charatan2024pixelsplat ; chen2025mvsplat ; xu2024depthsplat . Lastly, for ACID, we use only the test dataset, amounting to 1,972 scenes. These datasets cover a diverse range of scenes, including indoor and outdoor real estate videos (RE10K), aerial views of natural landscapes (ACID), and various real-world video scenes (DL3DV). We use an image resolution of 256 $\times$ 256 for the RE10K and ACID datasets, while for the DL3DV dataset, we use a resolution of 256 $\times$ 448 and 512 $\times$ 960.
@@ 4.2 Implementation and Training Details
Our model consists of 12 update layers, each containing one cross-attention block and one self-attention block. Inside each attention module, we adopt a pre-normalization method with LayerNorm ba2016layernorm and QK-Norm henry2020qknorm method with an RMSNorm zhang2019rmsnorm layer. Also, each of them utilizes multi-head attention vaswani2017attention with 12 heads and two linear layers with GELU hendrycks2016gelu activation. We set the hidden dimension of every linear layer to $d$ = 768, and use a patch size of $p$ = 8.
To enhance training efficiency in terms of both memory consumption and computational speed, we integrate several optimization techniques. Specifically, we utilize FlashAttention-2 dao2023flash-v2 to improve attention computation efficiency, apply gradient checkpointing chen2016checkpointing to reduce memory overhead, and adopt mixed-precision training with BFloat16 to accelerate computations while maintaining numerical stability. During training, we sample a wide range of input viewpoints based on the number of frames, and when necessary, apply farthest point sampling (FPS) qi2017pointnet++ using their camera positions. For each training example across all datasets, we sample 6 target views. For training on the DL3DV dataset, we start from the pretrained model on the RE10K dataset, following the approach in xu2024depthsplat . For more details, please refer to the supplementary material.
Figure 5: Qualitative comparison of novel view synthesis on the RE10K dataset.
Figure 6: Qualitative comparison of novel view synthesis on the DL3DV dataset.
@@ 4.3 Evaluation
We compare our model with recent generalizable 3D reconstruction methods charatan2024pixelsplat ; chen2025mvsplat ; zhang2025gs-lrm ; xu2024depthsplat ; nam2024generative . For evaluation, we follow the evaluation settings from charatan2024pixelsplat ; ye2024no for RE10K and xu2024depthsplat for DL3DV. Additionally, when utilizing more viewpoints than the evaluation protocol, we sample extra viewpoints between the input viewpoints, ensuring that these samples do not overlap with the target indices. We denote our various viewpoint settings as $(V,H/F,F)$, where $V$ is the number of viewpoints, and the following entries indicate the resolutions of viewpoint and image tokens (H: half-resolution, F: full-resolution). For example, a setting of $(2,H,F)$ indicates two viewpoints with half-resolution viewpoint tokens and full-resolution image tokens. $MC$ refers to our quarter mini-batch cross-attention (Fig. 4-(c)). Note that our 2-view full-resolution setting $(2,F,F)$ does not include token uplifting, as the viewpoint and image resolutions are identical.
4.4 Quantitative Results
We report quantitative results for the RE10K dataset in Tab. 1, for the DL3DV dataset in Tab. 2, and cross-dataset generalization results in Tab. 3. Our method consistently outperforms the baselines across various viewpoint configurations, including inference speed and memory usage, while achieving efficient scene representation with fewer Gaussians. These improvements are also reflected in the qualitative results shown in Fig. 5 and Fig. 6, where our method produces sharper view synthesis compared to the baselines, which exhibit noticeable artifacts. Note that, for the RE10K dataset, we improved the PSNR by 3$\sim$ 3∼ 3 dB over the state-of-the-art methods by leveraging more views (8 views vs. 2 views) with only less than half of the computation time (0.028s vs. 0.065s). For the DL3DV dataset, we improved the PSNR by 4$\sim$ 4∼ 4 dB over the state-of-the-art methods under the comparable computational budget, efficiently leveraging four times more views (6 vs. 24 views).

### v2

@@ Abstract
Feed-forward 3D modeling has emerged as a promising approach for rapid and high-quality 3D reconstruction. In particular, directly generating explicit 3D representations, such as 3D Gaussian splatting, has attracted significant attention due to its fast and high-quality rendering, as well as numerous applications. However, many state-of-the-art methods, primarily based on transformer architectures, suffer from severe scalability issues because they rely on full attention across image tokens from multiple input views, resulting in prohibitive computational costs as the number of views or image resolution increases. Toward a scalable and efficient feed-forward 3D reconstruction, we introduce an iterative Large 3D Reconstruction Model (iLRM) that generates 3D Gaussian representations through an iterative refinement mechanism, guided by three core principles: (1) decoupling the scene representation from input-view images to enable compact 3D representations; (2) decomposing fully-attentional multi-view interactions into a two-stage attention scheme to reduce computational costs; and (3) injecting high-resolution information at every layer to achieve high-fidelity reconstruction. Experimental results on widely used datasets, such as RE10K and DL3DV, demonstrate that iLRM outperforms existing methods in both reconstruction quality and speed.
@@ 1 Introduction
Since the recent success of 3D Gaussian Splatting (3D-GS) (Kerbl et al., 2023), significant progress has been made in leveraging this 3D representation for building generalizable feed-forward 3D reconstruction models (Charatan et al., 2024; Tang et al., 2025; Xu et al., 2024b; Chen et al., 2025; 2024d; Xu et al., 2025; Zhang et al., 2025). These methods typically train large neural networks to transform multi-view input images into feature representations, then regress Gaussian attributes. In contrast to per-scene 3D-GS optimization approaches (Kerbl et al., 2023; Mallick et al., 2024; Lu et al., 2024; Fang & Wang, 2024), these feed-forward models can reconstruct 3D scenes in a single forward pass, offering near real-time performance. Moreover, the prior knowledge learned from large-scale datasets (Zhou et al., 2018; Ling et al., 2024; Deitke et al., 2023; 2024) allows them to effectively generalize to unseen scenes. While their reconstruction quality often lags behind that of per-scene optimization methods, fast reconstruction speed and generalization capability mark a promising step toward the long-standing goal of achieving accurate and real-time 3D scene reconstruction.
@@
Among the promising approaches, pixel-aligned Gaussian models (Charatan et al., 2024; Szymanowicz et al., 2024b; Zheng et al., 2024) have emerged as the de facto standard, leveraging decades of advances in network architectures developed for numerous image-based tasks. While these models have proven effective, they also exhibit certain limitations. In particular, since they generate per-pixel Gaussians directly from the input images, the image resolution determines the number of Gaussians produced, which can lead to an excessive number of redundant Gaussians. For example, given input images at 1K resolution across 200 viewpoints (a scale comparable to the bicycle scene in the mip-NeRF 360 dataset (Barron et al., 2022)), these methods would produce 200 million Gaussians, despite prior studies (Lee et al., 2024; Fan et al., 2024; Chen et al., 2024c; Lee et al., 2025) demonstrating that such scenes can be efficiently represented with around 0.5 million Gaussians. To mitigate this issue, several techniques have been proposed, such as Gaussian regularization (Ziwen et al., 2025) and feature fusion (Wang et al., 2025b). Alternatively, the network architecture can also be designed to generate fewer Gaussians, for example by downsampling the output resolutions. However, these strategies still require processing high-resolution multi-view images and therefore do not address another fundamental limitation of these models: high computational and memory demands.
@@
A significant portion of computational and memory overhead arises from modeling interactions across multiple input views in feed-forward 3D reconstruction models. For instance, GS-LRM (Zhang et al., 2025) performs full attention over all image tokens from every input view, leading to a quadratic increase in complexity with respect to both the number of views and image resolution. MVSplat (Chen et al., 2025) and DepthSplat (Xu et al., 2025) construct and process dense cost volumes for each view, further contributing to the overall computational demands. While one might attempt to alleviate this burden by reducing the input image resolution or using a sparser set of views, such strategies risk discarding essential geometric and appearance information required for accurate reconstruction.
@@
Beyond the computational complexity and the inefficiency of the generated representations, we also question whether the prevailing formulation, casting 3D reconstruction as a sequence-to-sequence problem that maps entire sets of image tokens to dense, pixel-aligned Gaussians, is fundamentally well-suited to the nature of the task. While this formulation has achieved impressive results (even without explicit 3D inductive biases (Zhang et al., 2025; Ye et al., 2024)), it remains primarily a one-shot 3D scene generation process. In contrast, the recent optimization-based methods (Kerbl et al., 2023; Mallick et al., 2024) follow a fundamentally different strategy: they treat reconstruction as an iterative refinement process, where each iteration involves rendering the current scene estimate, measuring reconstruction error, and updating the representation accordingly. This loop enables the model to progressively capture finer geometric and appearance details while ensuring strong 3D consistency. The success of these methods suggests that high-quality 3D reconstruction may benefit not only from expressive representations but also from feedback-driven iterative refinement, a trait largely absent in existing feed-forward 3D models.
@@
We have comprehensively evaluated the proposed method on the large-scale datasets, RealEstate10K (Zhou et al., 2018) and DL3DV (Ling et al., 2024). The experimental results demonstrate that iLRM achieves superior rendering quality while substantially reducing both computational and memory overhead compared to recently proposed feed-forward Gaussian models. Moreover, in wide-coverage settings, our method outperforms optimization-based 3D-GS approaches, requiring only 0.5 seconds for inference while achieving higher fidelity reconstructions.
@@ 2.1 Motivation and Problem Statement
Existing generalizable 3D Gaussian reconstruction methods process multi-view images in an end-to-end fashion, often employing techniques such as epipolar line sampling (Charatan et al., 2024), plane-sweep stereo (Chen et al., 2025; 2024d; Xu et al., 2025), or full-resolution attention (Zhang et al., 2025; Xu et al., 2024b; Ziwen et al., 2025) to enforce multi-view consistency. While effective, these strategies introduce significant computational and memory overhead, limiting their scalability.
@@
We start with $N$ multi-view images $\{I_{i}\}_{i=1}^{N}$ and corresponding camera poses $\{C_{i}\}_{i=1}^{N}$. Based on this setup, our goal is to train a model $f_{\theta}$ that maps a set of viewpoints to 3D Gaussians, leveraging the associated multi-view images as visual cues throughout the reconstruction pipeline. More formally,
@@ 2.2 Architectural Design
Viewpoint tokenization. Following previous works (Tang et al., 2025; Zhang et al., 2025; Jin et al., 2024), we employ a Plücker ray embedding for each input view using the camera poses. Specifically, given the intrinsic, rotation, and translation, we construct the Plücker ray embeddings for each viewpoint. We then divide these viewpoint embeddings into non-overlapping patches of size $p\times p$, and reshape each patch into a 1D vector, resulting in a tensor of shape $H^{v}W^{v}/p^{2}\times 6p^{2}$. Then, we encode it using a single linear layer to produce viewpoint tokens, $V_{i}^{(0)}\in\mathbb{R}^{H^{v}W^{v}/p^{2}\times d}$. Plücker coordinates inherently capture spatial variations across pixels and views, allowing them to effectively differentiate between patches. As a result, we do not utilize additional positional embeddings.
@@
Figure 2: The proposed scalable architectural designs by decoupling viewpoint and image tokens, and modeling the global interactions via cross- and self-attentions (NN: # views, $h=H/p,w=W/p$).
Scalable multi-view context modeling. Fig. 2-(a) shows the typical feed-forward 3D methods (Zhang et al., 2025; Chen et al., 2025; Xu et al., 2025) using transformer architecture, which perform full attention across multi-view images, incurring a quadratic increase in computational cost with respect to both the number of views and the image resolution. Fig. 2-(b) depicts our decoupling approach. Thanks to the decoupling technique, we can reduce the viewpoint resolution while still leveraging high-resolution multi-view images (Fig. 2-(c)). We further decrease the computation cost by two-stage multi-view context modeling, per-view cross-attention and viewpoint self-attention (Fig. 2-(d)). For example, given 16 input images with a resolution of 256 $\times$ 256 and a patch size of 8, the relative computational cost follows the ratio (1:1:0.25:0.08, Fig. 2-(a):(b):(c):(d)), highlighting that our proposed method can accommodate more input views with significantly less computational burden.
@@
Figure 4: Various mini-batch cross-attention schemes.
@@ 2.3 Training Objectives
After generating 3D Gaussians from viewpoint tokens, we rasterize them to the target viewpoint to obtain rendered images, $\hat{I}_{t}$. These rendered images are then supervised using ground-truth images, $I_{t}$, through MSE loss and perceptual loss (Chen & Koltun, 2017; Li et al., 2020). For each training scene, our training loss function is as follows.
@@ 3.1 Datasets
We train our model on two large-scale datasets: RealEstate10K (RE10K) (Zhou et al., 2018), DL3DV (Ling et al., 2024), and evaluate it on three datasets, including ACID (Liu et al., 2021), which is used only for evaluation. We adopt the RE10K split following prior work (Charatan et al., 2024) and use the official split for DL3DV. We use an image resolution of 256 $\times$ 256 for the RE10K and ACID datasets, while for the DL3DV dataset, we use a resolution of 256 $\times$ 448 and 512 $\times$ 960. In addition, we employ the undistorted version of the DL3DV dataset at a resolution of 540 × 960, which originates from LongLRM (Ziwen et al., 2025).
@@ 3.2 Implementation and Training Details
Our model consists of 12 update layers, each containing one cross-attention and one self-attention block. Inside each attention module, we adopt a pre-normalization method with LayerNorm (Ba et al., 2016) and QK-Norm (Henry et al., 2020) method with an RMSNorm (Zhang & Sennrich, 2019) layer. Also, each of them utilizes multi-head attention (Vaswani et al., 2017) with 12 heads and two linear layers with GELU (Hendrycks & Gimpel, 2016) activation. We set the hidden dimension of every linear layer to $d$ = 768, and use a patch size of $p$ = 8.
To enhance training efficiency, we utilize FlashAttention-2 (Dao, 2023) to improve attention computation efficiency, apply gradient checkpointing (Chen et al., 2016) to reduce memory overhead, and adopt mixed-precision training with BFloat16 to accelerate computations while maintaining numerical stability. For more details, please refer to the supplementary material.
@@ 3.3 Evaluation
We compare our model against recent generalizable 3D reconstruction methods (Charatan et al., 2024; Chen et al., 2025; Zhang et al., 2025; Xu et al., 2025; Nam et al., 2025; Ziwen et al., 2025) as well as optimization-based approaches (Kerbl et al., 2023; Yu et al., 2024). For evaluation, we follow the evaluation settings from (Charatan et al., 2024; Ye et al., 2024) for RE10K and (Xu et al., 2025; Ziwen et al., 2025) for DL3DV. We denote our various viewpoint settings as $(V,H/F,F)$, where $V$ is the number of viewpoints, and the following entries indicate the resolutions of viewpoint and image tokens (H: half-resolution, F: full-resolution). For example, a setting of $(2,H,F)$ indicates two viewpoints with half-resolution viewpoint tokens and full-resolution image tokens. $MC$ refers to our quarter mini-batch cross-attention (Fig. 4-(c)). Note that our 2-view full-resolution setting $(2,F,F)$ does not include token uplifting, as the viewpoint and image resolutions are identical. Additionally, when utilizing more viewpoints than the evaluation protocol, we sample extra viewpoints between the input viewpoints, ensuring that these samples do not overlap with the target indices.
3.4 Results
In Tab. 12 and Fig. 5, we compare our approach with feed-forward methods on the RE10K dataset and cross-dataset generalization on ACID and DL3DV, which serve as the standard in related works. Furthermore, we report results with an increased number of input views (4 and 8), which incur less than half of the computation time compared to the baseline (DepthSplat) while achieving superior performance. For the DL3DV dataset, our method consistently outperforms the baselines across various viewpoint and resolution configurations, including inference speed and memory usage, while achieving efficient scene representation with fewer Gaussians, as shown in Tab. 34 and Fig. 67. While DepthSplat and our method are both trained under varying numbers of input views, our approach demonstrates enhanced scalability with respect to the number of views in Tab. 3.
In the wide-coverage setting (Tab. 5), following the LongLRM protocol (Ziwen et al., 2025), we evaluate using 32 high-resolution input images under full-frame coverage. For comparison, we also include optimization-based methods such as 3D-GS (Kerbl et al., 2023) and Mip-Splatting (Yu et al., 2024), both trained for 30k iterations using only the input images. LongLRM10 means finetuning 10 epochs initialized from the LongLRM’s generated Gaussians. Since our approach produces more compact 3D Gaussian representations, the finetuning process is significantly faster than the baseline.